# Zero-Shot Neural Architecture Search with Weighted Response Correlation

### v1

@@ 1 Introduction
The success of deep learning in various fields [1], especially computer vision, causes a surge in demand for designing neural architectures. Designing neural architectures manually requires extensive expertise and time investment. Neural architecture search (NAS) [1] offers a potential solution for automatically designing neural architectures, eliminating the need for human involvement.
@@
According to zero-shot NAS methods [13, 14, 15, 16, 17, 18] and potential proxies [21, 22, 23, 24, 25] from other fields, the performances of neural architectures mainly depend on their expressivity and generalizability. Based on this conclusion, we propose a novel training-free estimation proxy called weighted response correlation (WRCor), which can compute the response (activation and gradient) correlation coefficient matrices of different inputs throughout different layers of the network and aggregate them with layer-wise weights to calculate the proxy score. We clarify that a higher proxy score means the responses of different inputs are more linearly independent in a given network, which is positively related to its better performance. For further improvement, we propose two voting proxies, SPW and SJW, using the majority voting rule. The experimental results on proxy evaluation across three datasets and two NAS benchmarks verify the superiority of our training-free estimation proxies over existing proxies. The experimental results on architecture search across three search spaces demonstrate that our zero-shot NAS algorithms with various search strategies and our training-free estimation proxies outperform existing NAS algorithms for image recognition.
@@ 2 Related Work
Training-free proxy is the key component of zero-shot NAS. The goal of training-free proxies is to estimate the performance scores of neural architectures without training. NASWOT [13] uses the number of linear regions of the input space to measure the expressivity of neural architectures. TE-NAS [20] ranks architectures by analyzing the spectrum of the neural tangent kernel and the number of linear regions of the input space, which imply trainability and expressivity, respectively. Many saliency metrics of pruning-at-initialization are extended for zero-shot NAS [16], such as SNIP [21], GraSP [22], SynFlow [23], and Fisher [24, 26]. These proxies rely on the gradient w.r.t the parameters of neural networks, which are different approximations of the Taylor expansions of neural networks. Zen-score (Zen) [17] approximates $\Phi$-score (Phi) against model-collapse in very deep networks and can directly assert the model complexity. EPE-NAS [14] proposes an improvement of Jacobian correlation (JacCor) called EPE, which involves both intra- and inter-class correlation. ZiCo [18] reveals that the network with a high training convergence speed and generalization capacity should have high absolute mean values and low standard deviation values for the gradient w.r.t the parameters across different samples.
@@
There are also some potential proxies from other fields for training-free architecture estimation. Euclidean norm of parameters (PNorm) and Euclidean norm of parameter gradients (PGNorm) are the simplest saliency metrics for parameter pruning. We use the sum of Euclidean norms of all parameters or parameter gradients for architecture estimation. Barlow Twins loss (BT) [27] can measure the cross-correlation matrix between the outputs of two identical networks fed with distorted versions of a sample. We use Barlow Twins loss multiplied by -1 as a training-free proxy. Bundle entropy (BndlEnt) [28] is proposed to measure the sample conflict in each layer. We use the average value of bundle entropies for all network layers as a training-free proxy. Layer-wise dynamical isometry (LyrDynIso) [25] is defined as the singular values of the Jacobian matrix $\bm{J}^{l-1,l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ of layer $l$ are concentrated near 1 for all layers, which guarantees faithful signal propagation in the network. We also use a more efficient proxy DynIso that takes the Jacobian matrix $\bm{J}^{0,L}$ of the network into account instead of the layer-wise Jacobian matrix. The Jacobian matrix can measure the change in network output when the input changes. We use $l_{1}$-norm of the Jacobian matrix as a training-free proxy NetSens.
@@
Although training-free proxies of zero-shot NAS can significantly accelerate the NAS process, they fail to consistently and robustly outperform two naive proxies [15, 19, 18], i.e., Params and FLOPs. Besides, some proxies have specific requirements for neural architectures, e.g., NASWOT [13], TENAS [20], and ZenNAS [17]. These undermine the effectiveness, stability, and generality of existing zero-shot NAS methods with training-free proxies. Motivated by these limitations, we analyze the behavior of neural networks and propose an efficient, robust, and generic training-free estimation proxy called WRCor that offers a unified form for measuring both layer-wise expressivity and generalizability.
@@ 3.2 Generalizability
The generalizability refers to the generalization performance of neural architectures after training via gradient descent, which is equally important as expressivity. Intuitively, it is believed that gradients that exhibit greater linear dependence facilitate convergence [29] because when gradients are more linearly dependent, they point in similar directions, potentially leading to a more straightforward descent towards a minimum in the loss landscape. However, if the gradients are consistently aligned across different inputs, this could potentially lead to issues such as poor generalization or convergence to suboptimal minima. In essence, independent gradient directions might encourage to exploration of a broader range of solutions during training, thus enhancing its generalizability. Similarly, we also propose utilizing the correlation coefficient of gradients (correlation of gradients, GCor) across different inputs for evaluating the generalizability of neural architectures. By collecting the $i$-th flattened gradients $\bm{g}^{n}_{i}$ of the $n$-th sample by backward propagation, we compute the correlation coefficient matrices $\bm{\Sigma}^{G}_{i}$ for gradients $\bm{G}_{i}=(\bm{g}_{i}^{1},\bm{g}_{i}^{2},\cdots,\bm{g}_{i}^{N})^{\top}$. The main-diagonal elements of the correlation coefficient matrix $\bm{\Sigma}^{G}_{i}$ are always 1. Based on the above conclusion, networks with better generalizability have a lower correlation of the gradients across different inputs, i.e., smaller non-main-diagonal elements in the correlation coefficient matrix $\bm{\Sigma}^{G}_{i}$.
@@ 3.3.1 Correlation of Responses
For a $N$-order symmetric matrix ($N\geq 2$) where all main-diagonal elements are 1 and all non-main-diagonal elements are $x\in[0,1]$, its determinant value equals $[1+(N-1)x](1-x)^{N-1}$. It monotonically decreases w.r.t $x$. It attains the maximum value of 1 iff $x=0$ and the minimum value of 0 iff $x=1$.
@@
$\det\left(\begin{bmatrix}1&x&\cdots&x\\ x&1&\cdots&x\\ \vdots&\vdots&\ddots&\vdots\\ x&x&\cdots&1\\ \end{bmatrix}\right)=$ $[1+(N-1)x]\det\left(\begin{bmatrix}1&1&\cdots&1\\ x&1&\cdots&x\\ \vdots&\vdots&\ddots&\vdots\\ x&x&\cdots&1\\ \end{bmatrix}\right)$ (2)
$=$ $[1+(N-1)x]\det\left(\begin{bmatrix}1&1&\cdots&1\\ 0&1-x&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1-x\\ \end{bmatrix}\right)$ (3)
@@
$S_{ACor/GCor}=\log(\det(\bm{\Sigma}^{A/G}))\approx$ $N_{a}^{N}\det\left(\begin{bmatrix}1&x&\cdots&x\\ x&1&\cdots&x\\ \vdots&\vdots&\ddots&\vdots\\ x&x&\cdots&1\\ \end{bmatrix}\right)$ (6)
@@ 3.3.2 Weighted Correlation of Responses
Furthermore, our observations indicate that state-of-the-art networks often fail to distinguish different samples based on bottom-level features and rely on top-level features for clear discrimination. These networks usually capture common bottom-level features and independent top-level features. For superior neural architectures, the correlation of top-level responses is lower than that of bottom-level responses across different inputs. Based on our observation and hypothesis, we allow non-strict linear independence in bottom-level responses and focus on the correlation of top-level responses.
Specifically, when aggregating the correlation coefficient matrices, we assign higher weights to the correlation coefficient matrices of top-level responses than those of bottom-level responses. Given an $L$-layer neural network, where the $l$-th layer has $N^{l}_{a}$ non-linear activation units, we summarize these matrices using element-wise weighted summation. Intuitively, we empirically assign exponentially increasing weights from the bottom to the top, i.e.,
@@
In summary, compared to existing training-free proxies, our RCor proxy has four differences and advantages: 1) ours can apply to any neural architecture; 2) ours can consider both the expressivity and generalizability; 3) ours can simplify the calculation process by unifying score forms and matrix aggregation; 4) ours can consider correlation differences in different layers using the layer-wise score with exponentially increasing weights from the bottom to the top.
3.3.3 Voting Proxies
Given that no single proxy can consistently and robustly outperform all other proxies across all scenarios [30], proxy voting for architecture estimation is standard practice [16, 30]. To improve the performance and stability of proxies, we propose a voting proxy, named SJW, which employs a majority voting rule based on three highly competitive proxies: SynFlow, JacCor, and WRCor. For comparison, we also propose two additional voting proxies. ZeroCost [16] is a voting proxy of SynFlow, JacCor, and SNIP; SPW is a voting proxy of SynFlow, PNorm, and WRCor, where PNorm is the most cost-effective proxy except for Params and FLOPs. We do not use NASWOT and ZiCo as the voting experts because they have similar measures of the expressivity and generalizability of neural architectures.
@@ 3.4 Search Strategies
     $model.arch\leftarrow$ $controller$.GenerateArchitecture()
@@
     $controller$.Update($controller.reward$)
@@
  return highest-accuracy model in $history$
@@
     add $model$ to right of $population$
     add $model$ to $history$
@@
        $candidate\leftarrow$ random element from $population$
        add $candidate$ to $sample$
@@
     $parent\leftarrow$ highest-accuracy model in $sample$
@@
     add $child$ to right of $population$
     add $child$ to $history$
     remove $dead$ from left of $population$
     discard $dead$
@@
  return highest-accuracy model in $history$
@@ 4.1.1 Search Space and Benchmark
We validate our training-free proxies and NAS algorithms on three search spaces and benchmarks. NAS-Bench-101 search space [31] contains 423k unique convolutional architectures. NAS-Bench-101 benchmark provides the evaluation results of three runs on the CIFAR-10 [32] dataset for its 423k architectures. NAS-Bench-201 search space [33] consists of 15,625 possible architectures in total. NAS-Bench-201 benchmark gives the evaluation results on the CIFAR-10 [32], CIFAR-100 [32], and ImageNet16-120 [34] datasets for its 15,625 architectures. MobileNetV2 search space [35, 17] is a popular open-domain search space. Each architecture consists of stacked MobileNet Blocks. In each block, we can search for the depth-wise expansion ratio (E), the kernel size (K), the stride (S), the number of layers (L), the number of channels (C), the number of bottleneck channels (B), and the resolution of input images (R). The details of our MobileNetV2 search space are shown in Table 1.
@@ 4.1.3 Proxy Evaluation
We conduct the proxy evaluation experiments to prove the effectiveness of our training-free proxies. We use the Spearman ranking correlation coefficient (Spearman $\rho$) to quantify how well a proxy ranks architectures compared to the ground-truth ranking through training. For efficiency, we use 1000 randomly sampled architectures from the NAS-Bench-201 search space to estimate the real Spearman $\rho$ values of all architectures on the CIFAR-10 dataset for different proxies. We conduct multiple experiments across various settings, including different datasets (CIFAR-10, CIFAR-100, and ImageNet16-120 datasets), batch sizes (32, 64, and 128), initializations (Kaiming uniform, Kaiming normal, and $N(0,1)$ initializations), benchmarks (NAS-Bench-101 and NAS-Bench-201), and the numbers of sampled architectures (1000, top-20%, and 15625). The default setting of our proxies is bold in our experiments. All experiments are conducted on a single Tesla V100 GPU.
@@ 4.2 Analysis of Proxy Evaluation
Table 2: The comparison of the real/estimated Spearman $\rho$ values for different proxies on the NAS-Bench-201 benchmark. $\Delta$: the deviation between the real and estimated Spearman $\rho$ values; %: the deviation ratio. ValAcc: the validation accuracy through standard training. ValACC is a real performance metric and thus is the theoretical upper bound.
Proxies PNorm PGNorm SNIP GraSP SynFlow Fisher JacCor
estimated 0.684 0.591 0.593 0.543 0.743 0.506 0.726
real 0.684 0.594 0.596 0.515 0.739 0.503 0.726
$\Delta$ $0$0.003 -0.003 0.028 0.004 0.003 0.000
% 0.00 -0.51 -0.50 5.44 0.54 0.60 0.00
Proxies EPE Phi Zen BT BndlEnt LyDynIso DynIso
estimated 0.693 0.703 0.257 0.510 0.199 0.633 0.662
real 0.694 0.723 0.237 0.472 0.188 0.632 0.674
$\Delta$0.001 -0.020 0.020 0.038 0.011 0.001 -0.012
% -0.14 -2.77 8.44 8.05 5.85 0.16 -1.78
Proxies NetSens NASWOT ZiCo Params FLOPs
estimated 0.686 0.779 - - -
real 0.685 0.780 0.800 0.750 0.730
$\Delta$ 0.001 -0.001 - - -
% 0.15 -0.13 - - -
Proxies ACor RCor WRCor ZeroCost SPW SJW ValAcc
estimated 0.787 0.803 0.811 0.775 0.789 0.851 0.842
real 0.789 0.803 0.812 0.750 0.780 0.799 0.811
$\Delta$0.002 $0$0.001 0.025 0.009 0.052 0.031
% -0.25 0.00 -0.12 3.33 1.15 6.51 3.82
@@
Through experiments, we find that for superior architectures, the correlations of responses in their top layers are greater than those in their bottom layers. Conversely, for inferior architectures, the correlations in their top layers are worse than those in their bottom layers. For example, for the best NAS-Bench-101 architecture, the correlation scores in its first and last layers are -685.21 and -506.27 respectively; for the worst NAS-Bench-101 architecture, the scores in its first and last layers are -538.55 and -1098.02 respectively. Based on these observations, we propose a layer-wise weighting score, i.e., WRCor.
@@
To find a superior weighting way, we experiment with different ways of weighting the correlation of responses, including no weighting (no), weighting linearly (linear), quadratically (quad), and exponentially (expb2t), with weights increasing either from the bottom to the top layers or vice versa (expt2b). As shown in Table 3, the expb2t weighting way can achieve a notable improvement compared to linear and quad weighting ways. WRCor with the expb2t weighting has 0.008 (1.00%) higher Spearman $\rho$ value than RCor without weighting. The expt2b weighting harms architecture ranking and has 0.003 (0.37%) lower Spearman $\rho$ value than RCor. Experimental results demonstrate that the correlations of responses in the top layers are more significant than those in the bottom layers. The expb2t weighting adopted by our WRCor proxy outperforms other weighting ways.
Table 3: The comparison of different weighting ways for correlation of responses on the NAS-Bench-201 benchmark. $\Delta$: the deviation between RCor and the proxy; %: the deviation ratio.
Proxies no (RCor) linear quad expb2t (WRCor) expt2b
$\rho$ 0.803 0.805 0.806 0.811 0.800
$\Delta$ 0.000 0.002 0.003 0.008 -0.003
% 0.00 0.25 0.37 1.00 -0.37
@@
Figure 1 demonstrates the relationships between different estimated proxy metrics and the real test accuracy. Figure 1(a) and 1(b) demonstrate that the skewed distributions of SynFlow and JacCor proxies are not suitable for ranking these architectures and are the main reason for the low Spearman correlation. As expected, Figure 1(c) and Figure 1(d) show that the Params and FLOPs proxies exhibit excellent performance within each score segment, but it does not perform as well across all score segments collectively. Figure 1(e) and 1(f) show Acor and Rcor have a stronger correlation with the real test accuracy. Specifically, with the help of the correlation of gradients, the distribution in Figure 1(f) is approximately consistent with that of Figure 1(h). As shown in Figure 1(g), WRCor exhibits a more compact distribution by weighting the correlation of responses, which demonstrates that WRCor has a higher Spearman ranking correlation with the real test accuracy.
@@
Figure 1: The visualization of different proxies and test accuracy by 1000 random NAS-Bench-201 architectures.

### v2

@@ 1 Introduction
The success of deep learning in various fields [1], especially computer vision, causes a surge in demand for designing neural architectures. Designing neural architectures manually requires extensive expertise and time investment. Neural architecture search (NAS) [1, 2, 3, 4, 5, 6, 7] offers a potential solution for automatically designing neural architectures across various domains, eliminating the need for human involvement.
@@
According to zero-shot NAS methods [16, 17, 18, 19, 20, 21] and potential proxies [24, 25, 26, 27, 28] from other fields, the performances of neural architectures mainly depend on their expressivity and generalizability. Based on this conclusion, we propose a novel training-free estimation proxy called weighted response correlation (WRCor), as shwon in Figure 1, which can compute the response (activation and gradient) correlation coefficient matrices of different inputs throughout different layers of the network and aggregate them with layer-wise weights to calculate the proxy score. We clarify that a higher proxy score means the responses of different inputs are more linearly independent in a given network, which is positively related to its better performance. For further improvement, we propose two voting proxies, SPW and SJW, using the majority voting rule. The experimental results on proxy evaluation across three datasets and two NAS benchmarks verify the superiority of our training-free estimation proxies over existing proxies. The experimental results on architecture search across three search spaces demonstrate that our zero-shot NAS algorithms with various search strategies and our training-free estimation proxies outperform existing NAS algorithms for image recognition.
Figure 1: The overview of our proposed training-free proxy.
@@ 2 Related Work
Training-free proxy is the key component of zero-shot NAS. The goal of training-free proxies is to estimate the performance scores of neural architectures without training. NASWOT [16] uses the number of linear regions of the input space to measure the expressivity of neural architectures. However, NASWOT ignore the trainability of architectures, which imply how effective a network can be optimized via gradient descent. A network can achieve high performance only if the function it can represent is complex enough and at the same time, it can be effectively trained by gradient descent. In addition to expressivity, TE-NAS [23] propose to analyze the spectrum of the neural tangent kernel for measuring the trainability of architectures. According to this assumption that an untrained network with low correlation between different data points, where data points of the same category are closer to each other, can easily learn to distinguish the two data points during training, EPE-NAS [17] also proposes an improvement of Jacobian correlation (JacCor) [16] called EPE, which involves both intra- and inter-class correlation. Beside, many saliency metrics of pruning-at-initialization are extended for zero-shot NAS [19], such as SNIP [24], GraSP [25], SynFlow [26], and Fisher [27, 29]. These proxies rely on the gradient w.r.t the parameters of neural networks, which are different approximations of the Taylor expansions of neural networks. Zen-score (Zen) [20] approximates $\Phi$-score (Phi) against model-collapse in very deep networks and can directly assert the model complexity. Although these proxies perform well, their approximate expressions limit their effectiveness. None of these zero-shot proxies can actually work consistently better than a naive proxy, i.e., Params and FLOPs. Based on theoretical analysis, ZiCo [21] reveals that the network with a high training convergence speed and generalization capacity should have high absolute mean values and low standard deviation values for the gradient w.r.t the parameters across different samples.
@@
There are also some potential proxies from other fields for training-free architecture estimation. Euclidean norm of parameters (PNorm) and Euclidean norm of parameter gradients (PGNorm) are the simplest saliency metrics for parameter pruning. We use the sum of Euclidean norms of all parameters or parameter gradients for architecture estimation. Barlow Twins loss (BT) [30] can measure the cross-correlation matrix between the outputs of two identical networks fed with distorted versions of a sample. We use Barlow Twins loss multiplied by -1 as a training-free proxy. Bundle entropy (BndlEnt) [31] is proposed to measure the sample conflict in each layer. We use the average value of bundle entropies for all network layers as a training-free proxy. Layer-wise dynamical isometry (LyrDynIso) [28] is defined as the singular values of the Jacobian matrix $\bm{J}^{l-1,l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ of layer $l$ are concentrated near 1 for all layers, which guarantees faithful signal propagation in the network. We also use a more efficient proxy DynIso that takes the Jacobian matrix $\bm{J}^{0,L}$ of the network into account instead of the layer-wise Jacobian matrix. The Jacobian matrix can measure the change in network output when the input changes. We use $l_{1}$-norm of the Jacobian matrix as a training-free proxy NetSens. Although these potential proxies performs well in some scenes, their effectiveness in all situations are not experimentally verified or theoretically analyzed.
@@
These above training-free proxies can significantly accelerate the NAS process because they can estimate the performances of architectures without training. Compared with conventional NAS methods, zero-shot NAS is more efficient NAS method. However, there are three disadvantages for these above training-free proxies: 1) they fail to consistently and robustly outperform two naive proxies [18, 22, 21], i.e., Params and FLOPs; 2) they cannot consistently and robustly outperform other existing proxies in all situations; 3) some proxies have specific requirements for neural architectures, e.g., NASWOT [16] and TENAS [23] can only estimate the performance of CNN architecture and ZenNAS [20] requires a CNN architecture with ReLU as the activation function. These undermine the effectiveness, stability, and generality of existing zero-shot NAS methods with training-free proxies. To solve these limitations, we analyze the behavior of neural networks and propose an efficient, robust, and generic training-free estimation proxy called WRCor that offers a unified form for measuring both layer-wise expressivity and generalizability.
@@ 3.2 Generalizability
The generalizability refers to the ability to perform well on unseen data after being trained via gradient descent, which is equally important as expressivity. Intuitively, it is believed that gradients that exhibit greater linear dependence facilitate convergence [32] because when gradients are more linearly dependent, they point in similar directions, potentially leading to a more straightforward descent towards a minimum in the loss landscape. However, if the gradients are consistently aligned across different inputs, this could potentially lead to issues such as poor generalization or convergence to suboptimal minima. In essence, independent gradient directions might encourage to exploration of a broader range of solutions during training, thus enhancing its generalizability. Similarly, we also propose utilizing the correlation coefficient of gradients (correlation of gradients, GCor) across different inputs for evaluating the generalizability of neural architectures. By collecting the $i$-th flattened gradients $\bm{g}^{n}_{i}$ of the $n$-th sample by backward propagation, we compute the correlation coefficient matrices $\bm{\Sigma}^{G}_{i}$ for gradients $\bm{G}_{i}=(\bm{g}_{i}^{1},\bm{g}_{i}^{2},\cdots,\bm{g}_{i}^{N})^{\top}$. The main-diagonal elements of the correlation coefficient matrix $\bm{\Sigma}^{G}_{i}$ are always 1. Based on the above conclusion, networks with better generalizability have a lower correlation of the gradients across different inputs, i.e., smaller non-main-diagonal elements in the correlation coefficient matrix $\bm{\Sigma}^{G}_{i}$.
@@ 3.3.1 Correlation of Responses
For a $N$-order symmetric matrix ($N\geq 2$) where all main-diagonal elements are 1 and all non-main-diagonal elements are $x\in[0,1]$, its determinant value equals $1+(N-1)x](1-x)^{N-1}$. It monotonically decreases w.r.t $x$. It attains the maximum value of 1 iff $x=0$ and the minimum value of 0 iff $x=1$.
@@
[…]
@@
[…]
@@ 3.3.2 Weighted Correlation of Responses
Furthermore, our observations indicate that state-of-the-art networks often fail to distinguish different samples based on bottom-level features and rely on top-level features for clear discrimination. These networks usually capture common bottom-level features and independent top-level features. For superior neural architectures, the correlation of top-level responses is lower than that of bottom-level responses across different inputs. Based on our observation and hypothesis, we allow non-strict linear independence in bottom-level responses and focus on the correlation of top-level responses. Specifically, when aggregating the correlation coefficient matrices, we assign higher weights to the correlation coefficient matrices of top-level responses than those of bottom-level responses for weighting them. According to the above conclusion, we expect the importance of the responses to significantly increase with the number of layers. Therefore, we choose to exponentially weight these matrices. Compared to linear weighting and quadratic weighting, it places more emphasis on top-level responses. Their importances or weights increase geometrically across successive layers.
Given an $L$-layer neural network, where the $l$-th layer has $N^{l}_{a}$ non-linear activation units, we summarize these matrices using element-wise weighted summation. We assign exponentially increasing weights from the bottom to the top, i.e.,
@@
In summary, compared to existing training-free proxies, our WRCor proxy has three differences and advantages: 1) ours can consider both the expressivity and generalizability; 2) ours can simplify the calculation process by unifying score forms and matrix aggregation; 3) ours can consider correlation differences in different layers using the layer-wise score with exponentially increasing weights from the bottom to the top.
3.3.3 Proxy Voting
Proxy voting refers to the process where multiple proxies collaborate using predefined aggregation methods to estimate the performance of neural architectures. Given that no single proxy can consistently and robustly outperform all other proxies across all scenarios [33], proxy voting for architecture estimation is standard practice [19, 33]. To improve the performance and stability of proxies, we propose a voting proxy, named SJW, which employs a majority voting rule based on three highly competitive proxies: SynFlow, JacCor, and WRCor. For comparison, we also propose two additional voting proxies. ZeroCost [19] is a voting proxy of SynFlow, JacCor, and SNIP; SPW is a voting proxy of SynFlow, PNorm, and WRCor, where PNorm is the most cost-effective proxy except for Params and FLOPs. We do not use NASWOT and ZiCo as the voting experts because they have similar measures of the expressivity and generalizability of neural architectures.
@@ 3.4 Search Strategies
  $model.arch\leftarrow$ controllercontroller.GenerateArchitecture()
@@
  controllercontroller.Update($controller.reward$)
@@
return highest-accuracy model in historyhistory
@@
  add modelmodel to right of populationpopulation
  add modelmodel to historyhistory
@@
   $candidate\leftarrow$ random element from populationpopulation
   add candidatecandidate to samplesample
@@
  $parent\leftarrow$ highest-accuracy model in samplesample
@@
  add childchild to right of populationpopulation
  add childchild to historyhistory
  remove deaddead from left of populationpopulation
  discard deaddead
@@
return highest-accuracy model in historyhistory
@@ 4.1.1 Search Space and Benchmark
We validate our training-free proxies and NAS algorithms on three search spaces and benchmarks.
NAS-Bench-101 search space [34] contains 423k unique convolutional architectures. The NAS-Bench-101 architectures are composed of a stem layer followed by stacks of cells. Each cell is stacked 3 times, followed by a downsampling layer. This pattern is repeated 3 times, followed by global average pooling and a final dense softmax layer. Cells are represented as directed acyclic graphs (DAGs) with $V\leq 7$ nodes, $L=3$ candidate operations ($3\times 3$ convolution, $1\times 1$ convolution, and $3\times 3$ max-pool), and $E\leq 9$ edges. NAS-Bench-101 benchmark provides the evaluation results of three runs on the CIFAR-10 [35] dataset for its 423k architectures.
NAS-Bench-201 search space [36] consists of 15,625 possible architectures in total. The NAS-Bench-201 architectures follow the similar design of its counterpart as used in the NAS-Bench-101 benchmark. Each cell is represented as a densely connected DAG with $V=4$ nodes. The DAG is obtained by assigning a direction from the $i$-th node to the $j$-th node ($i<j$) for each edge in an undirected complete graph. Different NAS-Bench-101, their edges are associated with one of $L=5$ candidate operations (zeroize, skip connection, $3\times 3$ convolution, $1\times 1$ convolution, and $3\times 3$ average-pool). NAS-Bench-201 benchmark gives the evaluation results on the CIFAR-10 [35], CIFAR-100 [35], and ImageNet16-120 [37] datasets for its 15,625 architectures.
MobileNetV2 search space [38, 20] is a popular open-domain search space. Each architecture consists of stacked MobileNet Blocks. In each block, we can search for the depth-wise expansion ratio (E), the kernel size (K), the stride (S), the number of layers (L), the number of channels (C), the number of bottleneck channels (B), and the resolution of input images (R). The details of our MobileNetV2 search space are shown in Table 1.
@@ 4.1.3 Proxy Evaluation
We conduct the proxy evaluation experiments to prove the effectiveness of our training-free proxies. We use the Spearman ranking correlation coefficient (Spearman $\rho$) to quantify how well a proxy ranks architectures compared to the ground-truth ranking through training. For efficiency, we use 1000 randomly sampled architectures from the NAS-Bench-201 search space to estimate the real Spearman $\rho$ values of all architectures on the CIFAR-10 dataset for different proxies. We conduct multiple experiments across various settings, including different datasets (CIFAR-10, CIFAR-100, and ImageNet16-120 datasets), batch sizes (32, 64, and 128), initializations (Kaiming uniform, Kaiming normal, and $N(0,1)$ initializations), and benchmarks (NAS-Bench-101 and NAS-Bench-201). The default setting of our proxies is bold in our experiments. All experiments are conducted on a single Tesla V100 GPU.
@@ 4.2 Analysis of Proxy Evaluation
Table 2: The comparison of the real/estimated Spearman $\rho$ values for different proxies on the CIFAR-10 dataset and the NAS-Bench-201 benchmark. The real and estimated Spearman $\rho$ values are calculated by all architectures and 1000 randomly sampled architectures of the NAS-Bench-201 benchmark, respectively. $\Delta$: the deviation between the real and estimated Spearman $\rho$ values; %: the deviation ratio. ValAcc: the validation accuracy through standard training. ValACC is a real performance metric and thus is the theoretical upper bound.
Proxies est. real $\Delta$ %
PNorm 0.684 0.684 0.000 0.00
PGNorm 0.591 0.594 -0.003 -0.51
SNIP 0.593 0.596 -0.003 -0.50
GraSP 0.543 0.515 0.028 5.44
SynFlow 0.743 0.739 0.004 0.54
Fisher 0.506 0.503 0.003 0.60
JacCor 0.726 0.726 0.000 0.00
EPE 0.693 0.694 -0.001 -0.14
Phi 0.703 0.723 -0.020 -2.77
Zen 0.257 0.237 0.020 8.44
BT 0.510 0.472 0.038 8.05
BndlEnt 0.199 0.188 0.011 5.85
LyDynIso 0.633 0.632 0.001 0.16
DynIso 0.662 0.674 -0.012 -1.78
NetSens 0.686 0.685 0.001 0.15
NASWOT 0.779 0.780 -0.001 -0.13
ZiCo - 0.800 - -
Params - 0.750 - -
FLOPs - 0.730 - -
ValAcc 0.842 0.811 0.031 3.82
ACor (Ours) 0.787 0.789 -0.002 -0.25
RCor (Ours) 0.803 0.803 0.000 0.00
WRCor (Ours) 0.811 0.812 -0.001 -0.12
ZeroCost 0.775 0.750 0.025 3.33
SPW (Ours) 0.789 0.780 0.009 1.15
SJW (Ours) 0.851 0.799 0.052 6.51
@@
Through experiments, we find that for superior architectures, the correlations of responses in their top layers are greater than those in their bottom layers. Conversely, for inferior architectures, the correlations in their top layers are worse than those in their bottom layers. For example, for the best NAS-Bench-101 architecture, the correlation scores in its first and final layers are -685.21 and -506.27 respectively; for the worst NAS-Bench-101 architecture, the scores in its first and final layers are -538.55 and -1098.02 respectively. The scores of their final layers are positively correlated with their accuracies, which is consistent with our conclusion. The scores of their first layers are negatively correlated with their accuracies. This is not a common phenomenon because it is an extreme case where we are comparing the best and worst architectures. Based on these observations, we propose a layer-wise weighting score, i.e., WRCor.
@@
To find a superior weighting way, we experiment with different ways of weighting the correlation of responses, including no weighting (no), weighting linearly (linear), quadratically (quad), and exponentially (expb2t), with weights increasing either from the bottom to the top layers or vice versa (expt2b). As shown in Table 3, the expb2t weighting way can achieve a notable improvement compared to linearly and quadratically weighting ways. WRCor with the expb2t weighting has 0.008 (1.00%) higher Spearman $\rho$ value than RCor without weighting. The expt2b weighting harms architecture ranking and has 0.003 (0.37%) lower Spearman $\rho$ value than RCor. Experimental results demonstrate that the correlations of responses in the top layers are more significant than those in the bottom layers. The expb2t weighting adopted by our WRCor proxy outperforms other weighting ways. Besides, the above case of best and worst NAS-Bench-101 architectures also indicate that we should pay more attention the correlations of responses in the top layers and pay less attention or even ignore thoes in the bottom layers. Therefore, we adopt exponential weighting.
Table 3: The comparison of different weighting ways for correlation of responses on the CIFAR-10 dataset and the NAS-Bench-201 benchmark. $\Delta$: the deviation between RCor and the proxy; %: the deviation ratio.
Proxies $\rho$ $\Delta$ %
no (RCor) 0.803 0.000 0.00
linear 0.805 0.002 0.25
quad 0.806 0.003 0.37
expb2t (WRCor) 0.811 0.008 1.00
expt2b 0.800 -0.003 -0.37
@@
Figure 2 demonstrates the relationships between different estimated proxy metrics and the real test accuracy. Figure 2(a) and 2(b) demonstrate that the skewed distributions of SynFlow and JacCor proxies are not suitable for ranking these architectures and are the main reason for the low Spearman correlation. As expected, Figure 2(c) and Figure 2(d) show that the Params and FLOPs proxies exhibit excellent performance within each score segment, but it does not perform as well across all score segments collectively. Figure 2(e) and 2(f) show Acor and Rcor have a stronger correlation with the real test accuracy. Specifically, with the help of the correlation of gradients, the distribution in Figure 2(f) is approximately consistent with that of Figure 2(h). As shown in Figure 2(g), WRCor exhibits a more compact distribution by weighting the correlation of responses, which demonstrates that WRCor has a higher Spearman ranking correlation with the real test accuracy.
@@
Figure 2: The visualization of different proxies and test accuracy on the CIFAR-10 dataset by 1000 random NAS-Bench-201 architectures.